# Ferrotoroidal order of the monolayer ice in martyite

### v1

@@ Abstract
The beauty of snowflakes, the fractal patterns with the hexagonal symmetry, reflects the crystal structure of ice. The honeycomb network of H2O molecules stacks along the perpendicular direction, composing the hexagonal crystalline ice (ice Ih phase). What can one see if one layer of ice is extracted and confined to a two-dimensional (2D) sheet? Martyite Zn3(V2O7)(OH)2$\cdot$2H2O, a mineral with the honeycomb lattice of H2O in the porous framework, is an ideal system to study such monolayer ice. Due to the geometrical frustration and 2D nature of the honeycomb lattice, H2O molecules are dynamically disordered at room temperature, and the most stable molecular arrangement is nontrivial. In this study, we reveal disorder-order transitions of H2O in martyite using the single-crystal x-ray diffraction (XRD) and molecular dynamics (MD) simulation. The XRD results visualize the formation of hydrogen-bonded toroidal H2O hexamers, leading to the electric ferrotoroidal order below 200 K. Combined with the MD simulations, we discuss the formation process of the H2O hexamers and how they compromise the molecular arrangement towards lower temperatures. Our results unveil the ground state of monolayer ice, a fundamental knowledge to understand the polymorphism of H2O.
Main text
Figure 1: Crystal structures of martyite. (a) Crystal structure of martyite at room temperature. Three favoured orientations of H2O are shown for the A and B sublattices of the honeycomb lattice. Dashed lines show the hydrogen bond. (b) AB-site interaction energy scheme for the 3-states model. The stable geometry is arranged from left to right. (c) Ordered structure of H2O molecules in martyite at 30 K. The dotted and solid lines show the HT and LT unit cells, respectively. Representative interatomic distances are also shown. A perspective view of one H2O hexamer is shown in the right panel. (d) Ferrotoroidal order of the monolayer ice in martyite. The blue and green arrows show electric dipole moments of H2O and electric toroidal dipole moments of H2O hexamers, respectively.
@@
The polymorphism of water is a central interest of science. Ice shows diverse forms in different environments, from fluttering snowflakes to dense ice inside the core of planets [1, 2, 3]. The natural ice on the Earth (ice Ih) is a hydrogen-bonded crystal with a stack of buckled honeycomb layers of H2O molecules. The local configuration satisfies the two-covalent and two-hydrogen bonds in the tetrahedral geometry, the so-called ice rules (Bernal–Fowler rules) [4]. Since there are multiple ways to satisfy the ice rules, the proton configurations in ice Ih are disordered with Pauling’s residual entropy [5], reminiscent of the frustrating nature of ice. On the delicate balance of intermolecular interactions, at least 20 phases of ice have been reported in various temperature and pressure ranges [6, 7, 8, 9]. Among them, ices XVI and XVII are porous phases obtained by removing guest molecules from clathrate hydrates [10, 11, 12].
@@
Confinement is another approach to realizing novel states of water [13, 14]. Interfacial water formed on a metal surface or graphene sheets has been studied [15, 16, 17, 18]. Their local structures are similar to ice Ih, satisfying the ice rule, although the periodicity of the substrate globally influences the symmetry. Confined water inside minerals [19, 20, 21, 22, 23], molecular-organic frameworks [24, 25], and carbon nanotube [26, 27] has also been of interest. Here, the intermolecular distance and the coordination geometry can be controlled, extending the diversity of water to be explored.
@@
In this study, we focus on a hydrated mineral, martyite Zn3(V2O7)(OH)2$\cdot$2H2O, with the trigonal crystal system (space group $P\overline{3}m1$) [28, 29]. Martyite has the quasi-2D porous framework, made of the block layers of ZnO4(OH)2 octahedra separated by the V2O7 pillars (Fig. 1(a)). The water of crystallization confined in the porous framework forms the honeycomb network of H2O, which is isostructural to the monolayer ice. Each H2O is coordinated by the hydroxy group of the framework and forms the hydrogen bonds alternately in the ±$\pm c$ directions. This hydrogen bond moderately fixes the oxygen atom of H2O, while, the proton configuration (molecular orientation) can change in the $ab$ plane. As a result, the H2O molecules in martyite are dynamically disordered with the three favoured orientations along the nearest-neighbour directions of the honeycomb network, similar to the plastic ice VII [30].
@@
Our central question is, how these H2O molecules are arranged at low temperatures. As a first approximation, considering the three orientations at the A and B sites of the honeycomb lattice, four types of AB-site configurations are expected (Fig. 1(b)). Here, the configurations are listed in increasing order of energy from left to right, considering the hydrogen bond and dipole-dipole interaction. Note that H2O is a polar molecule with a local electric dipole moment pointing from the O atom toward the midpoint between the two H atoms. We comment that the out-of-plane interaction energy is of the order of 1 % of the in-plane interaction (see supplemental materials, SM [31]). Thus, water in martyite can be regarded as monolayer ice (quasi-2D system), where the AB-site interaction with the hydrogen bond plays a dominant role. On the honeycomb lattice of H2O, the average number of the hydrogen atoms per bonds is 4/3. Namely, all the nearest-neighbour molecules cannot be hydrogen-bonded. Therefore, the orientational degree of freedom of the monolayer ice is geometrically frustrated, leading to a highly degenerate ground state. In this study, we reveal the ground state of monolayer ice and discuss how the system lifts the frustration at low temperatures.
Dielectric properties
Figure 2: Dielectric anomalies of martyite. Temperature dependence of the dielectric constant $\varepsilon^{\prime}$ and dissipation factor $D$ of the polycrystalline martyite.
@@
Figure 2 shows the dielectric constant $\varepsilon^{\prime}$ and dissipation factor $D$ of the polycrystalline martyite as a function of temperature $T$. Two anomalies are observed at $T_{\mathrm{c1}}\approx$ 170–200 K and $T_{\mathrm{c2}}\approx$ 30–50 K. The decreases of $\varepsilon^{\prime}$ indicate the ordering of the H2O molecules rotating at room temperature. Based on the x-ray diffraction (XRD), Raman spectra, and molecular dynamics (MD) simulation, we conclude that these anomalies are related to the two structural phase transitions driven by the disorder-order transition of H2O (see following sections and SM [31]). Accordingly, we term the phases as HT ($T_{\mathrm{c1}}<T$), LT ($T_{\mathrm{c2}}<T<T_{\mathrm{c1}}$), and LT’ ($T<T_{\mathrm{c2}}$). The frequency dependences of $T_{\mathrm{c1}}$ and $T_{\mathrm{c2}}$ indicate the dielectric relaxation where the orientational degree of freedom gradually freezes. The observed dynamics reflects the energy barrier for rotational motions of H2O to relax to nearly-degenerate but lower energy states.
Single crystal XRD
Figure 3: Superlattice Bragg reflections of martyite. (a-d) XRD patterns of the single crystal of martyite at (a, b) 300 K and (c, d) 30 K. (e) Temperature dependence of the superlattice Bragg peak intensity at the index of $-4/3$, $5/3$, 0. (f) Temperature dependence of the lattice constants $a$ and $c$ of the original unit cell.
@@
Figures 3(a-d) show the single-crystal XRD patterns of martyite at 300 K and 30 K. We analyzed the XRD data at 300 K and confirmed that the crystal structure is consistent with the previous report [28, 29] (see SM [31]). Between the layers of martyite, the O atoms of H2O molecules form a slightly buckled honeycomb network with equivalent O-O bonds. Each O atom is bonded to three H atoms with a 2/3 occupancy, allowing the H2O molecule to retain the orientation degree of freedom of three-state vector (Fig. 1(a)). Superlattice peaks appear at the wave vector $q=\{1/3,1/3,0\}$ below $T_{\mathrm{c1}}\approx 200$ K, indicating a structural phase transition. Figure 3(e) shows the temperature dependence of the superlattice peak intensity -4/3 5/3 0 as an order parameter. No peak splitting or extinction rules are observed in any of the fundamental and superlattice peaks. Therefore, the unit cell becomes $\sqrt{3}a\times\sqrt{3}a\times c$, and the trigonal symmetry is maintained. We note that no superlattice peaks or diffuse scattering are observed along the $c^{*}$-axis in the low-temperature phase. This indicates that the ordering of H2O molecules is in-phase along the $c$ axis and is long-range ordered in martyite, in contrast to the proton-disordered ice Ih [32].
@@
We conclude that the space group of the LT phase at 30 K is $P\overline{3}$ (for details, see SM [31]). This lowering of symmetry generates merohedral twins due to the disappearance of mirror planes perpendicular to the tertiary axes of the trigonal lattice. In the current crystal, these domains exist in an approximately 1:1 ratio (see SM [31]). Figure 1(c) shows the ordered arrangement of H2O molecules at 30 K. The O atoms of the H2O molecules, which form a honeycomb lattice in the HT phase, are distorted into a breathing honeycomb lattice, forming H2O hexamers in the LT phase. In this hexamer, H2O molecules form a toroidal ring. One H atom in the H2O molecule points toward a neighbouring H2O molecule, while the other H atom is offset from the honeycomb plane. The off-plane H atoms are arranged alternately up and down as coordinated from the ZnO4(OH)2 block layer (Fig. 1(c)). This alternation in the monolayer ice resembles the hydrogen bonds in ice Ih formed in the ±$\pm c$ directions.
@@
Figure 3(f) shows the temperature dependence of the lattice constants. There is no anomaly in the $a$ and $c$ parameters, suggesting that the framework of martyite is rigid against the breathing deformation of the honeycomb network of H2O. The observed superlattice peak intensities are bilinear to the atomic displacements of O of H2O (Fig. S2(f) [31]).
MD simulation
@@
Figure 4: H2O hexamer and octadecamer formation in the MD simulation. (a) The number of hydrogen bonds per water molecule (blue) and the ratio of the type-1 H2O hexamers (red) as a function of temperature. The ratio of the type-1 hexamers is calculated for the designed ferrotoroidal state. (b) The in-plane radial distribution function between O atoms of H2O. The curves are shifted for clarity. (c) The snapshots of the hydrogen bonding network. The arrows represent the hydrogen-bonded pair defined by the intermolecular O–O distance ($d_{\mathrm{OO}}$) and the hydrogen bonding angle ($\theta$) in the right figure. Clockwise (counterclockwise) hexamers are marked by green crosses (yellow circles). (d) Spatial distribution of the hexamers, type 0, 1, and 2, simulated for the designed ferrotoroidal state. As shown in the right figures, purple (0), red (1), and blue (2) hexagons represent the H2O hexamers with zero, one, and two longer hydrogen bonds. The black dots indicate that at least one hydrogen bond in the hexamer is broken. The dotted triangles indicate the H2O octadecamers. The dashed line indicates the domain boundary of the lattice of the octadecamers.
@@
MD simulations on the monolayer ice in martyite demonstrate how the randomly oriented water molecules start to form hydrogen bonds with their neighbours. The simulation details and movie are given in SM [31]. In the simulation, we use the water model TIP4P/ICE [33] and design one layer of martyite, namely, the monlayer ice sandwitched by the ZnO4(OH)2 layers supported by the V2O7 pillars. All the structural parameters are taken from the XRD results. Water molecules interact with each other and with atoms of the framework via Lennard-Jones (LJ) and Coulomb potentials. Although no artificial constraints are imposed on the water molecules, their translational motions are naturally restricted to those around a honeycomb lattice site even at 300 K. No quantum effect on the dynamics of atoms is assumed in the MD simulations.
@@
Figure 4(a) shows the number of hydrogen bonds per water molecule $n_{\rm HB}$ as a function of temperature cooled down from 300 K. The results obtained with the TIP4P/2005 model are presented in SM [31], which are essentially the same as those with the TIP4P/ICE model. At 300 K, $n_{\rm HB}<1$, meaning that each water molecule in martyite assumes an almost random orientation. As the temperature decreases, $n_{\rm HB}$ gradually increases from 220 K, rises sharply between 150 and 200 K and reaches its maximum value of 2 at 100 K. Figure 4(c) depict the representative patterns of hydrogen bonds shown by the arrows at 80 K, 130 K, 180 K, and 300 K. At 300 K, the hydrogen-bonded molecules are sparse, among which the population of dimers is the largest. At the temperature of 180 K, where $n_{\rm HB}$ changes most rapidly, water molecules have one or two hydrogen bonds and together form structures of hexagonal rings and a variety of chains. At lower temperatures, water molecules form a regular array of hydrogen-bonded hexagonal rings. Each hexagonal ring has clockwise (CW) or counterclockwise (CCW) proton ordering as shown by the green crosses and yellow circles, respectively. The formation and breaking of hydrogen bonds takes place at high temperatures, but once the regular array of hexagonal rings is established, the hydrogen bonds are stable and no rearrangement occurs. The simulated transition from the orientationally disordered state (HT) to the array of hexagonal rings (LT) agrees with the XRD results. The transition temperature ($T_{\mathrm{c1}}$) also coincides with the experiment, indicating that the water molecules in martyite are well described by the monolayer ice model.
@@
In the H2O hexamer, H2O molecules form the toroidal ring structure. Here, a head-to-tail configuration of the electric dipole moments ($\mathbf{D}$) form an electric toroidal dipole moment defined by $\mathbf{A}=\sum\mathbf{r}\times\mathbf{D}$ (Fig. 1(d)). Our XRD results reveal that the LT phase of martyite shows an electric ferrotoroidal order driven by the disorder-order transition of water. So far, several compounds are known to show the electric ferrotoroidal order (also referred to as ferroaxial or ferro-rotational order, RbFe(MoO4)2, Ca5Ir3O12, and NiTiO3) [34, 35, 36, 37], which are driven by the rotational displacements of the coordination polyhedra. Amongst them, martyite is the simplest case where electric dipoles of polar molecules cooperatively form the toroidal motif, spontaneously breaking the mirror planes parallel to the $c$ axis. Here, the intermolecular interactions are relatively simple, and the MD simulation can reproduce the toroidal ordering. In the simulation, the signs of the toroidal moments (±$\pm\mathbf{A}$) are more or less random since the MD simulation is not long enough at each temperature (10 ns), and therefore the cooling rate is too fast to reach fully equilibrated states. However, once such an initial configuration is given by design, the ferrotoroidal order remains at least up to 100 K, and the internal energy is consistently lower than that of the simulated system without ferrotoroidal order (see SM [31]). The stability of the ferrotoroidal order can also be explained from the AB-site interaction energy scheme (Fig. 1(b) and see also Fig. S6 [31]). Thus, the ground state of the monolayer ice is a ferrotoroidal order, although the simulation gives the random toroidal pattern as a metastable state.
@@
To reveal the fine structure of the ground state, we design the forced ferrotoroidal state (all hexamers are CW) and see how the molecular arrangements change towards lower temperatures. Figures 4(d) show the snapshots of the water hexamers cooled down from the ferrotoroidal state. Here, type 0, 1, and 2 hexagons represent the H2O hexamers with zero, one, and two longer hydrogen bonds ($d_{\mathrm{OO}}>3.35$). At 120 K, various types of hexamers, including the broken ones shown by the black dots, are distributed randomly. By decreasing the temperature, the ratio of type-1 hexamers rapidly increases at $T_{\mathrm{c2}}\approx 100$ K (Fig. 4(a)). Here, three type-1 hexamers form a H2O octadecamer (dotted triangles in Fig. 4(d)). At the lowest temperature (LT’), H2O octadecamers alternatively align along one direction, leading to the orthorhombic distortion from the LT phase.
@@
The hexamers’ deformations are likely caused by the lattice constants of the honeycomb lattice, which are determined by the framework size. Specifically, the distance of 3.36 Å is significantly larger than 2.76 Å, the ideal hydrogen bond length in ice Ih. Thus, the H2O hexamer deforms for the sake of five stronger hydrogen bonds at the expense of one weaker one. Such structural instability is detected by the dielectric and Raman-scattering experiments at around $T_{\mathrm{c2}}\approx$ 30–50 K (Fig. 2 and Fig. S9 [31]).
@@
The MD simulation suggests that the LT’ phase is partially disordered (see the domain boundary shown by the dashed line in Fig. 4(d)). The energy cost to create a domain boundary is almost zero within our calculation accuracy. Besides, the distribution of the deformed hexamers depends on the water model and system size (see SM [31]). Therefore, the ground state of the martyite is highly degenerate due to the multiple ways of deforming the hexamers, indicating that the experimental realization of the long-range ordered H2O octadecamers might be challenging. Our XRD reveals that the atomic displacement parameters of O are elongated toward the intra-hexamer bond direction (see SM [31]), suggesting that local distortion from the regular hexamer may already occur at 30 K. However, the symmetry lowering from trigonal to orthorhombic structure is not detected within the experimental resolution, probably because 30 K is too close to $T_{\mathrm{c2}}$. Lower-temperature diffraction experiments are needed to discuss the symmetry of the LT’ phase.
Last, we comment on the slight discrepancy of $T_{\mathrm{c1}}$ observed in the dielectric (170–200 K) and XRD (200 K) experiments. The primary order parameter of the HT-LT phase transition is the electric toroidal dipole moment $\mathbf{A}$, which is not a polar multipole. Therefore, the dielectric measurement does not necessarily detect this ordering. The dielectric anomalies observed at $T_{\mathrm{c1}}$ (Fig. 2) probably reflect the ferrotoroidal domain dynamics, since the domain boundaries can locally possess electric dipole moments. Similarly, the deformed hexamers below $T_{\mathrm{c2}}$ possess electric dipole moments, and the alignment of the type-1 hexamers near the domain boundaries can be influenced by external electric fields. The frequency dependences of these dielectric anomalies ($T_{\mathrm{c1}}$ and $T_{\mathrm{c2}}$) indicate the energy barrier to move these domain boundaries. Therefore, the thermodynamical transition temperatures are expected to locate at the beginning of the dielectric relaxations, namely, $T_{\mathrm{c1}}\approx 200$ K and $T_{\mathrm{c2}}\approx 50$ K.

### v2

# Ferroaxial order of the monolayer ice in martyite

@@ Abstract
Ice Ih, the most stable phase of water at ambient pressure, is a stacking of the honeycomb network of water molecules H2O. What if one layer of ice is exfoliated and confined to a two-dimensional (2D) sheet? Martyite Zn3(V2O7)(OH)2$\cdot$2H2O, a mineral with the honeycomb lattice of H2O in the porous framework, is an ideal system for studying such monolayer ice. Due to the geometrical frustration and 2D nature, H2O molecules are dynamically disordered at room temperature. In this study, we reveal disorder-order transitions of H2O in martyite using single-crystal x-ray diffraction (XRD). The XRD results visualize the formation of hydrogen-bonded toroidal H2O hexamers, leading to the ferroaxial order below 200 K. Combined with the molecular dynamics simulations, we discuss the formation process of the H2O hexamers and how they compromise the molecular arrangement towards lower temperatures. Our results unveil the ground state of monolayer ice, a fundamental knowledge to understand the polymorphism of H2O.
@@
The polymorphism of water is a central interest of science. Ice shows diverse forms in different environments, from fluttering snowflakes to dense ice inside the core of planets [1, 2, 3]. The natural ice on the Earth (ice Ih) is a hydrogen-bonded crystal with a stack of buckled honeycomb layers of H2O molecules. The local configuration satisfies the two-covalent and two-hydrogen bonds in the tetrahedral geometry, the so-called ice rules (Bernal–Fowler rules) [4]. The ice rules allow for multiple proton configurations, causing disorder and residual entropy [5] in ice Ih and other ice forms, which manifests as an orientational disorder. Driven by a balance of intermolecular interactions, pressure-volume work, and entropy (including residual entropy), at least 20 ice phases have been reported across various temperature and pressure ranges [6, 7, 8, 9, 10, 11, 12].
@@
Confining water molecules offers a key strategy for accessing novel water states [13, 14], including the emergence of new kinds of phase transitions like liquid-solid critical points [15]. This approach has been used to study interfacial water on metal surfaces and graphene sheets [16, 17, 18, 19], where local structures resemble ice Ih (satisfying the ice rule), though the substrate’s periodicity globally influences the symmetry. Confinement within minerals [20, 21, 22, 23, 24, 25], molecular-organic frameworks [26, 27, 28], and carbon nanotubes [29, 30, 31] also allows for control over intermolecular distance and coordination geometry, expanding the diversity of water structures and phase behavior.
@@
In this study, we focus on a hydrated mineral, martyite Zn3(V2O7)(OH)2$\cdot$2H2O, with the trigonal crystal system (space group $P\overline{3}m1$) [32, 33]. Martyite has the quasi-2D porous framework, made of the block layers of ZnO4(OH)2 octahedra separated by the V2O7 pillars (Fig. 1). The quasi-2D water of crystallization forms the honeycomb network of H2O, which is isostructural to the monolayer ice. Each H2O is coordinated by the hydroxy group of the framework and forms the hydrogen bonds alternately in the ±$\pm c$ directions. This hydrogen bond moderately fixes the oxygen atom of H2O, while, the proton configuration (molecular orientation) can change in the $ab$ plane. As a result, the H2O molecules in martyite are dynamically disordered with the three favored orientations along the nearest-neighbor directions of the honeycomb network, similar to the plastic ice VII [34].
Figure 1: Crystal structure of martyite at room temperature. Three favored orientations of H2O are shown for the A and B sublattices of the honeycomb lattice. Dashed lines show the hydrogen bond.
@@
Our central question is, how these H2O molecules are arranged at low temperatures. As a first approximation, considering the three orientations at the A and B sites of the honeycomb lattice, four types of AB-site configurations are expected, and only one of them can form hydrogen bond (see supplemental materials, SM [35]). We comment that the out-of-plane interaction energy is of the order of 1 % of the in-plane interaction (see SM [35]). Thus, water in martyite can be regarded as monolayer ice (quasi-2D system), where the in-plane hydrogen bonding plays a dominant role. On the honeycomb lattice of H2O, the average number of hydrogen atoms per bond is 4/3. Namely, all the nearest-neighbor molecules cannot be hydrogen-bonded. Therefore, the orientational degree of freedom of the monolayer ice is geometrically frustrated, leading to a highly degenerate ground state. In this study, we reveal the ground state of monolayer ice and discuss how the system lifts the frustration at low temperatures.
We prepared single- and polycrystals of martyite by hydrothermal synthesis. We measured the dielectric properties of polycrystalline martyite pellets by using an LCR meter. We performed the x-ray diffraction (XRD) experiments using a single crystal in the temperature range from 30 to 300 K on BL02B1 at a synchrotron facility SPring-8 [36]. We carried out molecular dynamics (MD) simulations of two-dimensional water in martyite using GROMACS 2018 software [37] under periodic boundary conditions in the two dimensions. Experimental details are described in SM [35].
@@
Figure 2 shows the dielectric constant $\varepsilon^{\prime}$ and dissipation factor $D$ of the polycrystalline martyite as a function of temperature $T$. Two anomalies are observed at $T_{\mathrm{c1}}\approx$ 170–200 K and $T_{\mathrm{c2}}\approx$ 30–50 K. The decreases of $\varepsilon^{\prime}$ indicate the ordering of the H2O molecules rotating at room temperature. Based on the XRD, Raman spectra, and MD simulation, we conclude that these anomalies are related to the two structural phase transitions driven by the disorder-order transition of H2O (see following sections and Raman spectra data in SM [35]). Accordingly, we term the phases as HT ($T_{\mathrm{c1}}<T$), LT ($T_{\mathrm{c2}}<T<T_{\mathrm{c1}}$), and LT’ ($T<T_{\mathrm{c2}}$). The frequency dependences of $T_{\mathrm{c1}}$ and $T_{\mathrm{c2}}$ indicate the dielectric relaxation where the orientational degree of freedom gradually freezes. The observed dynamics reflect the energy barrier for rotational motions of H2O to relax to nearly-degenerate but lower energy states.
Figure 2: Temperature dependence of the dielectric constant $\varepsilon^{\prime}$ and dissipation factor $D$ of the polycrystalline martyite.
@@
Figures 3(a-d) show the single-crystal XRD patterns of martyite at 300 K and 30 K. We analyzed the XRD data at 300 K and confirmed that the crystal structure is consistent with the previous reports [32, 33] (see SM [35]). Between the layers of martyite, the O atoms of H2O molecules form a slightly buckled honeycomb network with equivalent O-O bonds. Each O atom is bonded to three H atoms with a 2/3 occupancy, allowing the H2O molecule to retain the orientation degree of freedom of the three-state vector (Fig. 1). Superlattice peaks appear at the wave vector $q=\{1/3,1/3,0\}$ below $T_{\mathrm{c1}}\approx 200$ K, indicating a structural phase transition. Figure 3(e) shows the temperature dependence of the superlattice peak intensity -4/3 5/3 0 as an order parameter. No peak splitting or extinction rules are observed in any of the fundamental and superlattice peaks. Therefore, the unit cell becomes $\sqrt{3}a\times\sqrt{3}a\times c$, and the trigonal symmetry is maintained. We note that no superlattice peaks or diffuse scattering are observed along the $c^{*}$-axis in the LT phase. This indicates that the ordering of H2O molecules is in-phase along the $c$ axis and is long-range ordered in martyite, in contrast to the proton-disordered ice Ih [38].
Figure 3: (a-d) XRD patterns of the single crystal of martyite at (a, c) 300 K and (b, d) 30 K. (e) Temperature dependence of the superlattice Bragg peak intensity at the index of $-4/3$, $5/3$, 0. (f) Temperature dependence of the lattice constants $a$ and $c$ of the original unit cell. (g) Ordered structure of H2O molecules in martyite at 30 K. The dotted and solid lines show the HT and LT unit cells, respectively. A perspective view of one H2O hexamer is shown in the panel. The blue and green arrows show electric dipole moments of H2O and electric toroidal dipole moments of H2O hexamers, respectively.
@@
We conclude that the space group of the LT phase at 30 K is $P\overline{3}$ (see SM [35]). This symmetry lowering generates merohedral twins due to the disappearance of mirror planes perpendicular to the tertiary axes of the trigonal lattice. In the current crystal, these domains exist in an approximately 1:1 ratio (see SM [35]). Figure 3(g) shows the ordered arrangement of H2O molecules at 30 K. The O atoms of the H2O molecules, which form a honeycomb lattice in the HT phase, are distorted into a breathing honeycomb lattice, forming H2O hexamers in the LT phase. In this hexamer, H2O molecules form a toroidal ring. One H atom in the H2O molecule points toward a neighboring H2O molecule, while the other H atom is offset from the honeycomb plane. The off-plane H atoms are arranged alternately up and down as coordinated from the ZnO4(OH)2 block layer. This alternation in the monolayer ice resembles the hydrogen bonds in ice Ih formed in the ±$\pm c$ directions.
@@
Figure 3(f) shows the temperature dependence of the lattice constants. There is no anomaly in the $a$ and $c$ parameters, suggesting that the framework of martyite is rigid against the breathing deformation of the honeycomb network of H2O. The observed superlattice peak intensities are bilinear to the atomic displacements of O of H2O (see SM [35]).
@@
Figure 4: (a) The number of hydrogen bonds per water molecule as a function of temperature. (b) The in-plane radial distribution function between O atoms of H2O. The curves are shifted for clarity. (c) The ratio of the hexamers formed on the honeycomb lattice. The black and red symbols represent all types of hexamers and type-1 hexamers, respectively. The ratio of the type-1 hexamers is calculated for the designed ferrotoroidal state. (d) The snapshots of the hydrogen bonding network. The arrows represent the hydrogen-bonded pair defined by the intermolecular O–O distance ($d_{\mathrm{OO}}$) and the hydrogen bonding angle ($\theta$) in the right figure. Clockwise (counterclockwise) hexamers are marked by green crosses (yellow circles). (e) Spatial distribution of the hexamers, type 0, 1, and 2, simulated for the designed ferrotoroidal state. As shown in the right figures, purple (0), red (1), and blue (2) hexagons represent the H2O hexamers with zero, one, and two longer hydrogen bonds. The black dots indicate that at least one hydrogen bond in the hexamer is broken. The dotted triangles indicate the H2O octadecamers. The dashed line indicates the domain boundary of the lattice of the octadecamers.
@@
MD simulations on the monolayer ice in martyite demonstrate how the randomly oriented water molecules start to form hydrogen bonds with their neighbors. The simulation movie is given in SM [35]. In the simulation, we use the water model TIP4P/ICE [39] and design one layer of martyite, namely, the monolayer ice sandwiched by the ZnO4(OH)2 layers supported by the V2O7 pillars, where the structural parameters are taken from the XRD results. Water molecules interact with each other and with atoms of the framework via Lennard-Jones and Coulomb potentials. Although no artificial constraints are imposed on the water molecules, their translational motions are naturally restricted to those around a honeycomb lattice site even at 300 K. No quantum effect on the dynamics of atoms is assumed in the MD simulations.
@@
Figure 4(a) shows the number of hydrogen bonds per water molecule $n_{\rm HB}$ as a function of temperature cooled down from 300 K. The results obtained with the TIP4P/2005 model are presented in SM [35], which are essentially the same as those with the TIP4P/ICE model. At 300 K, $n_{\rm HB}<1$, meaning that each water molecule in martyite assumes an almost random orientation. As the temperature decreases, $n_{\rm HB}$ gradually increases from 220 K, rises sharply between 140 and 200 K, and reaches its maximum value of 2 at 100 K. Figure 4(d) depict the representative patterns of hydrogen bonds shown by the arrows at 80 K, 130 K, 180 K, and 300 K. At 300 K, the hydrogen-bonded molecules are sparse, among which the population of dimers is the largest. At the temperature of 180 K, where $n_{\rm HB}$ changes most rapidly, water molecules have one or two hydrogen bonds and together form structures of hexagonal rings and a variety of chains. Below $T_{\mathrm{c1}}\approx 150$ K, the number of hydrogen-bonded hexamers increase rapidly (Fig. 4(c)), forming a regular array of hexamers. Each hexamer has clockwise (CW) or counterclockwise (CCW) proton ordering as shown by the green crosses and yellow circles, respectively. The formation and breaking of hydrogen bonds takes place at high temperatures, but once the regular array of hexamers is established, the hydrogen bonds are stable and no rearrangement occurs. The simulated transition from the orientationally disordered state (HT) to the array of hexamers (LT) agrees with the XRD results, indicating that the water molecules in martyite are well described by the monolayer ice model.
@@
In the H2O hexamer, H2O molecules form the toroidal ring structure. Here, a head-to-tail configuration of the electric dipole moments ($\mathbf{D}$) form an electric toroidal dipole moment defined by $\mathbf{A}=\sum\mathbf{r}\times\mathbf{D}$ (Fig. 3(g)). Our XRD results reveal that the LT phase of martyite is a ferrotoroidal state of H2O driven by the disorder-order transition. Such ordering is referred to as a ferroaxial or ferro-rotational order, usually driven by the rotational displacements of the coordination polyhedra (e.g. RbFe(MoO4)2, Ca5Ir3O12, NiTiO3, and 1T-TaS2) [40, 41, 42, 43, 44]. Among them, martyite is the simplest case where electric dipoles of polar molecules cooperatively form the toroidal motif, spontaneously breaking the mirror planes parallel to the $c$ axis. Here, the intermolecular interactions are relatively simple, and the MD simulation can reproduce the toroidal ordering. In the simulation, the signs of the electric toroidal moments (±$\pm\mathbf{A}$) are more or less random since the MD simulation is not long enough at each temperature (10 ns), and therefore the cooling rate is too fast to reach fully equilibrated states. However, once such an initial configuration is given by design, the ferrotoroidal order remains at least up to 100 K, and the internal energy is consistently lower than that of the simulated system without ferrotoroidal order (see SM [35]). Thus, the ground state of the monolayer ice is a ferrotoroidal order, although the simulation gives the random toroidal pattern as a metastable state.
@@
To reveal the fine structure of the ground state, we design the forced ferrotoroidal state (all hexamers are CW) and see how the molecular arrangements change towards lower temperatures (Fig. 4(e)). Here, type 0, 1, and 2 hexagons represent the H2O hexamers with zero, one, and two longer hydrogen bonds ($d_{\mathrm{OO}}>3.35$). At 120 K, various types of hexamers, including the broken ones shown by the black dots, are distributed randomly. By decreasing the temperature, the ratio of type-1 hexamers rapidly increases at $T_{\mathrm{c2}}\approx 100$ K (Fig. 4(c)). Here, three type-1 hexamers form a H2O octadecamer (dotted triangles in Fig. 4(e)). At the lowest temperature (LT’), H2O octadecamers alternatively align along one direction, leading to the orthorhombic distortion from the LT phase.
@@
The hexamers’ deformations are likely caused by the lattice constants of the honeycomb lattice, which are determined by the framework size. Specifically, the distance of 3.36 Å is significantly larger than 2.76 Å, the ideal hydrogen bond length in ice Ih. Thus, the H2O hexamer deforms for the sake of five stronger hydrogen bonds at the expense of one weaker one. Such structural instability is detected by the dielectric and Raman-scattering experiments at $T_{\mathrm{c2}}\approx$ 30–50 K (Fig. 2 and Fig. S7 [35]).
@@
The MD simulation suggests that the LT’ phase is partially disordered (see the domain boundary shown by the dashed line in Fig. 4(e)). The energy cost to create a domain boundary is almost zero within our calculation accuracy. Besides, the distribution of the deformed hexamers depends on the water model and system size (see SM [35]). Therefore, the ground state of the martyite is highly degenerate due to the multiple ways of deforming the hexamers, indicating that the experimental realization of the long-range ordered H2O octadecamers might be challenging. Our XRD reveals that the atomic displacement parameters of O are elongated toward the intra-hexamer bond direction (see SM [35]), suggesting that local distortion from the regular hexamer may already occur at 30 K. However, the symmetry lowering from trigonal to orthorhombic structure is not detected within the experimental resolution, probably because 30 K is too close to $T_{\mathrm{c2}}$. Lower-temperature diffraction experiments are needed to discuss the symmetry of the LT’ phase.